# Posterior Ramifications of Prior Dependence Structures

### v2

@@ Abstract
In fully Bayesian analyses, prior distributions are specified before observing data. Prior elicitation methods transfigure prior information into quantifiable prior distributions. Recently, methods that leverage copulas have been proposed to accommodate more flexible dependence structures when eliciting multivariate priors. We prove that under broad conditions, the posterior cannot retain many of these flexible prior dependence structures in large-sample settings. We emphasize the impact of this result by overviewing several objectives for prior specification to help practitioners select prior dependence structures that align with their objectives for posterior analysis. Because correctly specifying the dependence structure a priori can be difficult, we consider how the choice of prior copula impacts the posterior distribution in terms of asymptotic convergence of the posterior mode. Our resulting recommendations streamline the prior elicitation process.
@@ 1.1 Overview of Prior Elicitation
Statistical methods leverage data to infer properties about an unobservable, possibly multivariate parameter $\boldsymbol{\theta}=(\theta_{1},...,\theta_{d})$. Bayesian methods for statistical inference (see e.g., Gelman et al. (2013)) require the specification of a prior distribution for the parameter $\boldsymbol{\theta}$, denoted by $p(\boldsymbol{\theta})$. This distribution characterizes the beliefs about $\boldsymbol{\theta}$ prior to observing any data. For a particular statistical model, $L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y})$ is the relevant likelihood function for the parameter $\boldsymbol{\theta}$ with respect to the observed data, denoted by the vector or matrix $\boldsymbol{y}$. Bayesian inference employs Bayes’ theorem to update the beliefs about the random variable $\boldsymbol{\theta}$ as more information becomes available via the observed data. In the Bayesian paradigm, inference is facilitated via the posterior distribution of $\boldsymbol{\theta}$, denoted by $p(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y})$. This distribution communicates which values of $\boldsymbol{\theta}$ are plausible given the observed data and prior beliefs. By Bayes’ Theorem, we have that
@@
Winkler (1967) conducted some of the initial work on prior elicitation, citing the siloed nature of prior specification and posterior analysis. Most previous Bayesian research had investigated how to leverage sample information from the likelihood function to obtain the posterior from the prior distribution, which was assumed to have been already assessed. Winkler (1967) explored elicitation methods for Bernoulli processes, two of which involved directly asking questions about the prior cumulative distribution function (CDF) or probability distribution function (PDF) for the Bernoulli parameter. These elicitation methods were for univariate priors, yet Winkler (1967) acknowledged that the assessment of prior distributions in multivariate contexts was an important and nontrivial problem.
@@
Decades later, Chaloner (1996) overviewed methods for subjective prior specification with an emphasis on methods for clinical trials. In this context, Freedman and Spiegelhalter (1983) elicited a prior on the difference in cancer recurrence probability between treatment groups by asking experts about the mode and likely lower and upper bounds for this difference. This prior was not combined with observed data and instead used for design purposes to choose the number of interim analyses in a sequential trial. Other contributions in clinical settings advocated for soliciting prior beliefs from several experts to form a community of prior distributions and basing inference on a consensus of posterior conclusions (Kadane, 1986; Chaloner et al., 1993). To reduce cognitive and computational complexity, most elicitation methods overviewed by Chaloner (1996) relied on parametric assumptions and solicited information about potentially observable conduits for the unobservable parameter $\boldsymbol{\theta}$ that are easier to conceptualize. For instance, priors for Bayesian regression model coefficients were elicited using information about quantiles of predictive distributions (Kadane et al., 1980) and survival probabilities (Chaloner et al., 1993).
@@
Additional reviews of prior elicitation methods have since been conducted. Garthwaite et al. (2005) divided the elicitation process into four stages: setup, elicitation, fitting, and adequacy. These stages involve preparing the expert for elicitation, eliciting distributional summaries of expert knowledge, fitting a probability distribution to these summaries, and assessing the adequacy of elicitation. O’Hagan et al. (2006) focused on elicitation of subjective probabilities and noted that experts tend to characterize events as impossible instead of assigning them small probabilities. Because this gives rise to hard boundaries in the support of the elicited priors, O’Hagan et al. (2006) overviewed methods to correct for expert overconfidence. Johnson et al. (2010) conducted a systematic review of prior elicitation methods with an emphasis on the feasibility of the process in terms of the required time, cost, personnel, and equipment. They found that although expert fatigue and lack of understanding compromise the reliability of elicitation procedures (Winkler, 1971; Garthwaite et al., 2005), the reviewed methods were rarely formally evaluated on their feasibility.
@@
Recent contributions have aimed to reduce expert fatigue and improve understanding by developing iterative elicitation procedures that provide experts with instant feedback via a graphical interface (Jones and Johnson, 2014; Casement and Kahle, 2018; Williams et al., 2021; Casement and Kahle, 2023). Prior elicitation procedures have also been developed for more complex settings – including methods for rank analysis (Crispino and Antoniano-Villalobos, 2023), nonparametric models (Seo and Kim, 2022), sequential analysis (Santos and Costa, 2019), power priors (Ye et al., 2022), and mixture models (Fúquene et al., 2019; Feroze and Aslam, 2021). One such area of research involves using copulas to accommodate more flexible dependence structures when eliciting multivariate priors (Elfadaly and Garthwaite, 2017; Wilson, 2018; Wilson et al., 2021). These contributions are novel given that the dependence structure is an afterthought in many recent prior elicitation methods. In particular, many recent methods consider elicitation for univariate parameters (Casement and Kahle, 2018, 2023), fix the dependence structure to leverage conjugacy (Santos and Costa, 2019; Srivastava et al., 2019), or assume that the components of $\boldsymbol{\theta}$ are independent a priori (Garthwaite et al., 2013; Jones and Johnson, 2014; Fúquene et al., 2019; Seo and Kim, 2022). Prominent recent advances with copula-based priors are overviewed in Section 1.3. Copula-based priors are the focus of this paper, and a primer on copulas is provided in Section 1.2.
@@ 1.2 Background on Copula Models
where $\Phi$ is the CDF of the $\mathcal{N}(0,1)$ distribution and $\Phi_{\boldsymbol{R}}:\mathbb{R}^{d}\rightarrow[0,1]$ is the CDF of the $d$-dimensional $\mathcal{N}(\boldsymbol{0},\boldsymbol{R})$ distribution with correlation and covariance matrix $\boldsymbol{R}$. The $t$-copula (Demarta and McNeil, 2005) can be similarly defined given a multivariate $t$-distribution with correlation matrix $\boldsymbol{R}$ and degrees of freedom $\nu$. In contrast, Archimedean copulas are a commonly used class of copulas that are efficiently parameterized via generator functions (Nelsen, 2006). This paper focuses on parametric copula models, but copulas can also be leveraged in a nonparametric framework (Wong and Ma, 2010; Wu et al., 2015; Ning and Shephard, 2018; Barone and Dalla Valle, 2023).
For $d=2$ dimensions, Figure 1 visualizes samples from two copulas. In the left plot, the blue points are generated from an Archimedean Clayton copula with parameter value $\phi=3\geq 0$, which characterizes positive dependence. The extent of the dependence between two random quantities is constrained by the Fréchet-Hoeffding bounds (Nelsen, 2006). The dotted line in the left plot of Figure 1 is the upper Fréchet-Hoeffding bound. This bound characterizes dependence for comonotonic variables. In the right plot, the blue points are generated from a Gaussian copula parameterized by Pearson’s $\rho=-0.8$. The dotted line in the right plot is the lower Fréchet-Hoeffding bound. This bound characterizes dependence for countermonotonic variables. As the strength of the positive (negative) dependence between two variables increases, their $(u_{1},u_{2})$ combinations tend to cluster around the upper (lower) Fréchet-Hoeffding bound. The dependence structure between two independent random variables is characterized by the independence copula: $c(\boldsymbol{u})=1$ for $\boldsymbol{u}\in[0,1]^{2}$. In two dimensions, a sample from this copula approximates random scatter over $[0,1]^{2}$.
Figure 1: Samples of 1000 points from a Clayton copula (left) and a Gaussian copula with Pearson’s $\rho=-0.8$ (right). The upper and lower Fréchet-Hoeffding bounds are given by the dotted lines.
The upper Fréchet-Hoeffding bound for positive dependence generalizes to settings with more than two dimensions. However, negative dependence in higher dimensions is more complicated since many random variables cannot exhibit strong mutual negative dependence. In addition to characterizing the strength and direction of dependence between random variables, copula models also account for symmetry and tail dependence. For instance, the Clayton copula accounts for lower tail dependence since the points in the left plot of Figure 1 are tightly clustered in the bottom left corner.
@@ 1.3 Recent Developments with Copula-Based Priors
Copulas can be incorporated into the prior elicitation process as illustrated in recent developments for the multinomial model. Elfadaly and Garthwaite (2017) proposed one such method to elicit Gaussian copula prior distributions. The standard multinomial model assumes that data $y_{i}\in\{1,2,\dots,w\},~{}i=1,\dots,n$ are collected independently and that the outcome $v$ occurs with probability $10<p_{v}<$ 1 for $v=1,\dots,w$ such that $\sum_{v=1}^{w}p_{v}=1$. The multinomial model is parameterized by $\boldsymbol{p}=(p_{1},\dots,p_{w})$. However, Elfadaly and Garthwaite (2017) did not directly assign a Gaussian copula to the multinomial probabilities since that approach would not enforce the unit-sum constraint. Instead, they defined new variables $Z_{1},\dots,Z_{w}$ such that
$\begin{split}Z_{1}=p_{1},~{}~{}~{}~{}Z_{v}=\dfrac{p_{v}}{1-\sum_{t=1}^{v-1}p_{t}}~{}~{}\text{for}~{}~{}v=2,\dots,w-1,~{}~{}~{}~{}\text{and}~{}~{}~{}~{}Z_{w}=1.\end{split}$ (2)
@@
The variable $Z_{v}$ represents the probability that an observation is assigned to category $v$ given that it has not been assigned to categories $11,\dots,v-$ 1. Elfadaly and Garthwaite (2017) assigned marginal $\text{BETA}(\alpha_{v},\beta_{v})$ priors to $Z_{v}$, $v=1,\dots,w-1$. A joint prior for $\boldsymbol{\theta}=\boldsymbol{Z}_{w-1}=(Z_{1},\dots,Z_{w-1})$ that satisfies the unit-sum constraint was created by joining the marginal beta priors with a Gaussian copula. If $Z_{1},\dots,Z_{w-1}$ are independent, $\boldsymbol{p}_{w}$ follows a generalized Dirichlet distribution (Connor and Mosimann, 1969). In this scenario, Gaussian copulas were leveraged to construct priors that were more flexible than standard alternatives.
Elfadaly and Garthwaite (2017) constructed marginal beta distributions for $Z_{v},v=1,\dots,w-1$ by soliciting estimates for the quartiles of each variable. For each variable, the three quartile estimates were reconciled into a two-parameter beta distribution using least-squares optimization. They formed the correlation matrix $\boldsymbol{R}$ for the Gaussian copula by soliciting further estimates. For $v=2,\dots,w-1$, experts were asked to update their estimate for the median of $p_{v}$ under the assumption that the median of $p_{v-1}$ was equal to the lower quartile specified in the previous step. These additional estimates were used in conjunction with a method from Kadane et al. (1980) to ensure $\boldsymbol{R}$ was positive definite.
@@
Wilson (2018) extended this method for use with vine copulas (Joe, 1996; Bedford and Cooke, 2002; Joe and Kurowicka, 2011), using a similar process as Elfadaly and Garthwaite (2017) to elicit marginal beta distributions for $Z_{v},v=1,\dots,w-1$. This extended method leveraged D-vines (Kurowicka and Cooke, 2005) to incorporate more flexibility into the prior dependence structure than the Gaussian copula can accommodate. For a model with $d$ variables, D-vines utilize the graphical structure in Figure 2 to characterize dependence between the variables in $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{d})$ using $d-1$ trees: $T_{1},\dots,T_{d-1}$. $T_{1}$ consists of a node set $N_{1}=\{1,2,\dots,d\}$ and an edge set $E_{1}=\{(1,2),(2,3),\dots,(d-1,d)\}$, where the integers in the node and edge sets refer to indices in $\boldsymbol{\theta}$. For $j=2,3,\dots,d-1$, the node set of $T_{j}$ is $N_{j}=E_{j-1}$, and two edges in $E_{j-1}$ are connected with an edge in $T_{j}$ only if they share a common node. D-vines characterize dependence in higher-dimensional settings using (un)conditional bivariate copulas. Each edge $\boldsymbol{e}$ in the edge set $E(\mathcal{V})=\cup_{j=1}^{d-1}E_{j}$ considers the dependence between two variables in $\boldsymbol{\theta}$, denoted $e_{1}$ and $e_{2}$, conditional on the variables in a potentially empty set $D_{\boldsymbol{e}}\subset\{1,2,\dots,d\}$.
Figure 2: The structure of a general D-Vine on $d$ variables.
Wilson (2018) proposed considering Gaussian and $t$-copulas as candidate copulas, along with several Archi-medean copulas that are absolutely continuous with full support. These bivariate copulas were selected by soliciting estimates for the conditional quartiles of the $p_{v}$ and $Z_{v}$ variables. Across all candidate copula families considered, the best-fitting copula was parameterized to minimize least squares between solicited and induced prior quantiles on the $Z_{v}$ variables.

### v3

@@ Abstract
Prior elicitation methods for Bayesian analyses transfigure prior information into quantifiable prior distributions. Recently, methods that leverage copulas have been proposed to accommodate more flexible dependence structures when eliciting multivariate priors. We show that the posterior cannot retain many of these flexible prior dependence structures in large-sample settings, and we emphasize that it is our responsibility as statisticians to communicate this to practitioners. We therefore overview objectives for prior specification that guide conversations between statisticians and practitioners to promote alignment between the flexibility in the prior dependence structure and the objectives for posterior analysis. Because correctly specifying the dependence structure a priori can be difficult, we consider how the choice of prior copula impacts the posterior distribution in terms of asymptotic convergence of the posterior mode. Our resulting recommendations clarify when it is useful to elicit intricate prior dependence structures and when it is not.
@@ 1.1 Overview of Prior Elicitation
Statistical methods leverage data to infer properties about an unobservable, possibly multivariate parameter $\boldsymbol{\theta}=(\theta_{1},...,\theta_{d})$. Bayesian methods for statistical inference (see e.g., (Gelman et al., 2013)) require the specification of a prior distribution for the parameter $\boldsymbol{\theta}$, denoted by $p(\boldsymbol{\theta})$. This distribution characterizes the beliefs about $\boldsymbol{\theta}$ prior to observing any data. For a particular statistical model, $L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y})$ is the likelihood function for the parameter $\boldsymbol{\theta}$ with respect to the observed data, denoted by the vector or matrix $\boldsymbol{y}$. Bayesian inference employs Bayes’ theorem to update the beliefs about the random variable $\boldsymbol{\theta}$ as more information becomes available via the observed data. In the Bayesian paradigm, inference is facilitated via the posterior distribution of $\boldsymbol{\theta}$, denoted by $p(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y})$. This distribution communicates which values of $\boldsymbol{\theta}$ are plausible given the observed data and prior beliefs. By Bayes’ Theorem, we have that
@@
Winkler (Winkler, 1967) conducted some of the initial work on prior elicitation, citing the siloed nature of prior specification and posterior analysis. Most previous Bayesian research had investigated how to leverage sample information from the likelihood function to obtain the posterior from the prior distribution, which was assumed to have been already assessed. Winkler (Winkler, 1967) explored elicitation methods for Bernoulli processes that involved asking questions about the prior cumulative distribution function (CDF) or probability distribution function (PDF) for the Bernoulli parameter. These elicitation methods were for univariate priors, yet Winkler (Winkler, 1967) acknowledged that the assessment of prior distributions in multivariate contexts was an important and nontrivial problem.
@@
Decades later, Chaloner (Chaloner, 1996) overviewed methods for subjective prior specification with an emphasis on methods for clinical trials. In this context, Freedman and Spiegelhalter (Freedman and Spiegelhalter, 1983) elicited a prior on the difference in cancer recurrence probability between treatment groups by asking experts about the mode and likely lower and upper bounds for this difference. This prior was not combined with observed data and instead used for design purposes to choose the number of interim analyses in a sequential trial. Other contributions in clinical settings advocated for soliciting prior beliefs from several experts to form a community of prior distributions and basing inference on a consensus of posterior conclusions (Kadane, 1986; Chaloner et al., 1993). To reduce cognitive and computational complexity, most elicitation methods overviewed by Chaloner (Chaloner, 1996) relied on parametric assumptions and solicited information about potentially observable conduits for the unobservable parameter $\boldsymbol{\theta}$ that are easier to conceptualize. For instance, priors for Bayesian regression model coefficients were elicited using information about quantiles of predictive distributions (Kadane et al., 1980) and survival probabilities (Chaloner et al., 1993).
@@
Additional reviews of prior elicitation methods have since been conducted (Garthwaite et al., 2005; O’Hagan et al., 2006; Johnson et al., 2010). In one contribution, Johnson et al. (Johnson et al., 2010) conducted a systematic review of prior elicitation methods with an emphasis on the feasibility of the process in terms of the required time, cost, personnel, and equipment. They found that although expert fatigue and lack of understanding compromise the reliability of elicitation procedures (Winkler, 1971; Garthwaite et al., 2005), the reviewed methods were rarely formally evaluated on their feasibility.
@@
Recent contributions have aimed to reduce expert fatigue and improve understanding by developing iterative elicitation procedures that provide experts with instant feedback via a graphical interface (Jones and Johnson, 2014; Casement and Kahle, 2018; Williams et al., 2021; Casement and Kahle, 2023). Prior elicitation procedures have also been developed for more complex settings – including methods for rank analysis (Crispino and Antoniano-Villalobos, 2023), nonparametric models (Seo and Kim, 2022), sequential analysis (Santos and Costa, 2019), power priors (Ye et al., 2022), and mixture models (Fúquene et al., 2019; Feroze and Aslam, 2021). One such area of research involves using copulas to accommodate more flexible dependence structures when eliciting multivariate priors (Elfadaly and Garthwaite, 2017; Wilson, 2018; Wilson et al., 2021). These contributions are novel given that the dependence structure is an afterthought in many recent prior elicitation methods. In particular, many recent methods consider elicitation for univariate parameters (Casement and Kahle, 2018, 2023), fix the dependence structure to leverage conjugacy (Santos and Costa, 2019; Srivastava et al., 2019), or assume that the components of $\boldsymbol{\theta}$ are independent a priori (Garthwaite et al., 2013; Jones and Johnson, 2014; Fúquene et al., 2019; Seo and Kim, 2022). Prominent recent advances with copula-based priors are overviewed in Section 1.3. A primer on copulas is provided in Section 1.2.
@@ 1.2 Background on Copula Models
where $\Phi$ is the CDF of the $\mathcal{N}(0,1)$ distribution and $\Phi_{\boldsymbol{R}}:\mathbb{R}^{d}\rightarrow[0,1]$ is the CDF of the $d$-dimensional $\mathcal{N}(\boldsymbol{0},\boldsymbol{R})$ distribution with covariance matrix equal to the correlation matrix $\boldsymbol{R}$. The $t$-copula (Demarta and McNeil, 2005) can be similarly defined given a multivariate $t$-distribution with correlation matrix $\boldsymbol{R}$ and degrees of freedom $\nu$. In contrast, Archimedean copulas are a commonly used class of copulas that are efficiently parameterized via generator functions (Nelsen, 2006). This paper focuses on parametric copula models, but copulas can also be leveraged in a nonparametric framework (Wong and Ma, 2010; Wu et al., 2015; Ning and Shephard, 2018; Barone and Dalla Valle, 2023).
For $d=2$ dimensions, Figure 1 visualizes samples from two copulas. In the top plot, the blue points are generated from an Archimedean Clayton copula with parameter value $\phi=3\geq 0$, which characterizes positive dependence. The extent of the dependence between two random quantities is constrained by the Fréchet-Hoeffding bounds (Nelsen, 2006). The dotted line in the top plot of Figure 1 is the upper Fréchet-Hoeffding bound. This bound characterizes dependence for comonotonic variables. In the bottom plot, the blue points are generated from a Gaussian copula parameterized by Pearson’s $\rho=-0.8$. The dotted line in the bottom plot is the lower Fréchet-Hoeffding bound. This bound characterizes dependence for countermonotonic variables. As the strength of the positive (negative) dependence between two variables increases, their $(u_{1},u_{2})$ combinations tend to cluster around the upper (lower) Fréchet-Hoeffding bound. The dependence structure between two independent random variables is characterized by the independence copula: $c(\boldsymbol{u})=1$ for $\boldsymbol{u}\in[0,1]^{2}$. In two dimensions, a sample from this copula approximates random scatter over $[0,1]^{2}$.
Figure 1: Samples of 1000 points from a Clayton copula (top) and a Gaussian copula with Pearson’s $\rho=-0.8$ (bottom). The upper and lower Fréchet-Hoeffding bounds are given by the dotted lines.
The upper Fréchet-Hoeffding bound for positive dependence generalizes to settings with more than two dimensions. However, negative dependence in higher dimensions is more complicated since many random variables cannot exhibit strong mutual negative dependence. In addition to characterizing the strength and direction of dependence between random variables, copula models also account for symmetry and tail dependence. For instance, the Clayton copula accounts for lower tail dependence since the points in the top plot of Figure 1 are tightly clustered in the bottom left corner.
@@ 1.3 Recent Developments with Copula-Based Priors
Copulas can be incorporated into the prior elicitation process as illustrated in recent developments for the multinomial model. Elfadaly and Garthwaite (Elfadaly and Garthwaite, 2017) proposed one such method to elicit Gaussian copula prior distributions. The standard multinomial model assumes that data $y_{i}\in\{1,2,\dots,w\},~{}i=1,\dots,n$ are collected independently and that the outcome $v$ occurs with probability $10<p_{v}<$ 1 for $v=1,\dots,w$ such that $\sum_{v=1}^{w}p_{v}=1$. The multinomial model is parameterized by $𝑤\boldsymbol{p}_{\color[rgb]{0,0,0}{w}}\color[rgb]{0,0,0}{=}(p_{1},\dots,p_{w})$. However, Elfadaly and Garthwaite (Elfadaly and Garthwaite, 2017) did not directly assign a Gaussian copula to the multinomial probabilities since that approach would not enforce the unit-sum constraint. Instead, they defined new variables $Z_{1},\dots,Z_{w}$ such that
$\begin{split}Z_{1}=p_{1},~{}~{}~{}~{}Z_{v}=\dfrac{p_{v}}{1-\sum_{t=1}^{v-1}p_{t}}~{}~{}\text{for}~{}~{}v=2,\dots,w-1,\\ \text{and}~{}~{}~{}~{}Z_{w}=1.\end{split}$ (2)
@@
Figure 2: The structure of a general D-Vine on $d$ variables. For each tree $T_{j}$, the node set $N_{j}$ inside the rectangles is joined by the edge set $E_{j}$, labelled using the text over the line segments. These labels take the form $e_{1},e_{2}\hskip 2.0pt|\hskip 2.0ptD_{\boldsymbol{e}}$.
The variable $Z_{v}$ represents the probability that an observation is assigned to category $v$ given that it has not been assigned to categories $11,\dots,v-$ 1. Elfadaly and Garthwaite (Elfadaly and Garthwaite, 2017) assigned marginal $\text{BETA}(\alpha_{v},\beta_{v})$ priors to $Z_{v}$, $v=1,\dots,w-1$. A joint prior for $\boldsymbol{\theta}=\boldsymbol{Z}_{w-1}=(Z_{1},\dots,Z_{w-1})$ that satisfies the unit-sum constraint for $\boldsymbol{p}_{w}$ was created by joining the marginal beta priors with a Gaussian copula. If $Z_{1},\dots,Z_{w-1}$ are independent, $\boldsymbol{p}_{w}$ follows a generalized Dirichlet distribution (Connor and Mosimann, 1969). In this scenario, Gaussian copulas were leveraged to construct priors that were more flexible than standard alternatives.
Elfadaly and Garthwaite (Elfadaly and Garthwaite, 2017) constructed marginal beta distributions for $Z_{v},v=1,\dots,w-1$ by soliciting estimates for the quartiles of each variable. For each variable, the three quartile estimates were reconciled into a two-parameter beta distribution using least-squares optimization. They formed the correlation matrix $\boldsymbol{R}$ for the Gaussian copula by soliciting further estimates. For $v=2,\dots,w-1$, experts were asked to update their estimate for the median of $p_{v}$ under the assumption that the median of $p_{v-1}$ was equal to the lower quartile specified in the previous step. These additional estimates were used in conjunction with a method from Kadane et al. (Kadane et al., 1980) to ensure $\boldsymbol{R}$ was positive definite.
@@
Wilson (Wilson, 2018) extended this method for use with vine copulas (Joe, 1996; Bedford and Cooke, 2002; Joe and Kurowicka, 2011), using a similar process as Elfadaly and Garthwaite (Elfadaly and Garthwaite, 2017) to elicit marginal beta distributions for $Z_{v},v=1,\dots,w-1$. This extended method leveraged D-vines (Kurowicka and Cooke, 2005) to incorporate more flexibility into the prior dependence structure than the Gaussian copula can accommodate. For a model with $d$ variables, D-vines utilize the graphical structure in Figure 2 to characterize dependence between the variables in $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{d})$ using $d-1$ trees: $T_{1},\dots,T_{d-1}$. $T_{1}$ consists of a node set $N_{1}=\{1,2,\dots,d\}$ and an edge set $E_{1}=\{(1,2),(2,3),\dots,(d-1,d)\}$, where the integers in the node and edge sets refer to indices in $\boldsymbol{\theta}$. For $j=2,3,\dots,d-1$, the node set of $T_{j}$ is $N_{j}=E_{j-1}$, and two edges in $E_{j-1}$ are connected with an edge in $T_{j}$ only if they share a common node in $N_{j-1}$. D-vines characterize dependence in higher-dimensional settings using (un)conditional bivariate copulas. Each edge $\boldsymbol{e}$ in the edge set $E(\mathcal{V})=\cup_{j=1}^{d-1}E_{j}$ considers the dependence between two variables in $\boldsymbol{\theta}$, denoted $e_{1}$ and $e_{2}$, conditional on the variables in a potentially empty set $D_{\boldsymbol{e}}\subset\{1,2,\dots,d\}$. For an edge $\boldsymbol{e}$, $D_{\boldsymbol{e}}$ consists of all indices that are common to both nodes connected by the edge, and the two remaining indices are $e_{1}$ and $e_{2}$.
The magnitude and direction of dependence in bivariate copulas is often summarized via Kendall’s $\tau$ (Kendall, 1938) values. Kendall’s $\tau$ measures rank correlation in terms of how similar the orderings of bivariate data are when ranked by each quantity. The D-vine structure allows the bivariate copula corresponding to each edge $\boldsymbol{e}\in E(\mathcal{V})$ to take any Kendall’s $\tau$ value that does not correspond to comonotonicity or countermonotonicity (Bedford and Cooke, 2002): $\tau_{e_{1},e_{2}|D_{\boldsymbol{e}}}\in(-1,1)$.
When used to specify a prior $p(\boldsymbol{\theta})$, D-vines prompt the set of Kendall’s $\tau$ values $\{\tau^{p}_{e_{1},e_{2}|D_{\boldsymbol{e}}}\}_{\boldsymbol{e}\in E(\mathcal{V})}$. For each bivariate prior copula, its family determines the range of possible values for Kendall’s $\tau$. Wilson (Wilson, 2018) proposed considering Gaussian and $t$-copulas as candidate copulas, along with several Archimedean copulas that are absolutely continuous with full support. These bivariate copulas were selected by soliciting estimates for the conditional quartiles of the $p_{v}$ and $Z_{v}$ variables. Across all candidate copula families considered, the best-fitting copula was parameterized to minimize least squares between solicited and induced prior quantiles on the $Z_{v}$ variables.